# Exploring the Privacy-Energy Consumption Tradeoff for Split Federated Learning

### 05 Jan 2024 v2

@@ 3.2 Other Types of Control Variables
  • Computation resource control: The computation resource (denoted as $f_{i}$) of each client participating in SFL plays a pivotal role in optimizing performance. However, due to the early stage of this field, there has been a lack of rigorous analysis for SFL. Fundamentally, $f_{i}$ is known to affect computation latency, denoted as $T_{i}$ in our analysis, which in turn impacts energy consumption within the field of FL [11].
  • Power control: As studied in other studies of FL [11], the clients participating in the SFL can also improve their performance by optimizing their transmission power $P_{t}$. This optimization has a significant impact on both uplink transmission rates, which were assumed to be constant in our initial analysis for simplicity, and energy consumption during the transmission period. Consequently, we can further enhance energy efficiency and convergence speeds by adjusting transmission power.
  • Radio resource management: Similar to power control, the allocation of radio resources (i.e., bandwidth allocation over clients) has been regarded as a crucial control parameter managed by base stations. This is because efficient bandwidth allocation can significantly impact both the uplink and downlink transmission rates of clients participating in FL. In the case of SFL, where more data transmission is required than FL process, such as client models and smashed data exchanges between clients and the main server or fed server, radio resource allocation can be designed with a more elaborate approach.
@@
  • Client selection: The high mobility of clients can affect the wireless network conditions between clients and the main server as well as the fed server, respectively. Furthermore, the clients often have unbalanced and non-independent and identically distributed (non-IID) datasets, which can result in slower convergence speeds. Therefore, taking these issues into account, it can be beneficial to explore dynamic client selection management within SFL to enhance both accuracy and convergence speed.
As we reviewed, the performance of SFL, concerning both energy efficiency and convergence speed, etc., is noteworthy. However, achieving an optimal balance with respect to these multiple variables often requires substantial complexity in the pursuit of a global optimum. In some instances, adopting a modular design approach can be advantageous, transforming the overall problem into several sub problems such as a convex optimization one. Even when dealing with non-convex scenarios, a range of heuristic techniques can be devised to efficiently attain sub-optimal solution.
4 Open Challenges
4.1 Deep Reinforcement Learning
@@
Given the dynamic nature of networking and the computing capabilities of clients, the deep reinforcement learning (DRL) holds the potential to strike a balance between privacy and energy consumption in the SFL by intelligently selecting the appropriate cut layer, in conjunction with other essential control parameters, as we have discussed. In this context, the key challenges in applying DRL to address various SFL issues revolve around precisely defining the agent, environment, state, action, and reward. Additionally, the selection of an appropriate DRL model and the fine-tuning of hyperparameters become pivotal for improving convergence speed, especially when dealing with complex and dynamic environments.
4.2 Privacy and Security Protection
As in [14], when clients and the main server hold their data and labels, respectively, it is important to protect these from exposure due to privacy concerns (e.g., vehicles communicating with roadside infrastructure, or smart remote sensing where network owners collaborate in training). In such situations, attackers could be malicious data owners or eavesdroppers trying to intercept gradients shared from the main server to clients. The similarity in cut layer gradients among data samples can disclose their connection to labels, which is known as a label inference attack. A common technique for this attack involves using clustering mechanisms, assuming the attacker is aware of the number of classes and has auxiliary data to initiate the clustering algorithm. Therefore, to mitigate risks, both reconstruction and label inference attacks must be addressed by carefully choosing the cut layer and implementing DP. Moreover, security attacks to SFL can significantly degrade model training, bringing it down to unacceptable performance levels due to actions by malicious clients. Given that SFL involves communication with both the fed server and the main server, it is more vulnerable to attacks. These attacks could involve directly poisoning the client-side local model or the smashed data. Therefore, it is crucial to explore advanced methods for detecting malicious clients as a defense mechanism to enhance the security of SFL.
4.3 Quantization Optimization
@@
To enhance the efficiency of SFL, it is beneficial to consider the quantization approach. This technique has gained extensive attention in the fields of FL [15] and holds promise for addressing the significant communication overhead observed between clients and servers. In the context of SFL, adopting a quantization approach can substantially reduce the size of both local model updates and intermediate data. This reduction leads to several advantages, including reduced bandwidth usage, lower transmission energy, and reduced storage needs. Furthermore, quantization introduces inherent privacy guarantees. This is because by reducing the precision of transmitted data, it inherently obfuscates the exact values, thereby adding an extra layer of protection for sensitive information, which is crucial in maintaining data privacy. Consequently, it is advisable to further research and develop an appropriate quantization approach tailored to SFL. This research direction aims to achieve the above benefits while maintaining an acceptable level of accuracy.
5 Conclusions

### 20 Mar 2024 v3

@@ III-B Other Types of Control Variables
  • Computation/Power/Radio resource management: The computation resource, transmission power, and radio resource (i.e., bandwidth allocation over clients) of each client participating in SFL plays a pivotal role in optimizing performance. However, due to the early stage of this field, there has been a lack of rigorous analysis for SFL. Fundamentally, those factors are known to affect overall latency, denoted as $T_{i}$ in our analysis, which in turn impacts energy consumption within the field of FL [10].
@@
  • Client and dataset management: The high mobility of clients can affect the wireless network conditions between clients and the main server and the fed server, respectively. Additionally, variations in the size of datasets across clients, characterized by their unbalanced, non-independent, and identically distributed (non-IID) nature, may lead to slower convergence speeds. Given these challenges, implementing strategies for dynamic client selection and effective dataset management—such as downsampling or augmenting datasets—within SFL can be beneficial in improving both accuracy and convergence speeds.
As we reviewed, the performance of SFL, concerning both energy efficiency and convergence speed, etc., is noteworthy. The heterogeneity in clients’ computing and networking capabilities can impact the speed of convergence due to the issue of stragglers, and this diversity also leads to differences in energy consumption among clients. Correspondingly, achieving an optimal balance concerning these multiple variables often requires substantial complexity in pursuing a global optimum. This necessitates adopting a modular design strategy or employing various heuristic methods to achieve sub-optimal solutions.
IV Open Challenges
IV-A Deep Reinforcement Learning
@@
Considering the dynamic and heterogeneous nature of client networking and computing capabilities, deep reinforcement learning (DRL) shows promise in intelligently selecting the optimal cut layer. This aims to balance privacy, energy consumption, and convergence speed in the SFL, along with other essential control parameters, as discussed. Recent research by [15] illustrated how DRL is utilized to optimize cut layer selection for collaborative DL inference in multi-access computing (MEC) servers by efficiently offloading certain inference tasks from clients to the MEC server. The key challenges in applying DRL to address various SFL issues revolve around precisely defining the agent, environment, state, action, and reward. Additionally, selecting an appropriate DRL model and fine-tuning hyperparameters become pivotal for improving convergence speed, especially when dealing with complex and dynamic environments.
IV-B Privacy and Security Protection
As in [9], when clients and the main server hold their data and labels, respectively, attackers could be malicious data owners or eavesdroppers trying to intercept gradients shared from the main server to clients. The similarity in cut layer gradients among data samples can disclose their connection to labels, known as a label inference attack. Therefore, to mitigate risks, label inference attacks must be addressed by carefully choosing the cut layer and implementing differential privacy (DP). Moreover, determining the cut layer is also critical for maintaining security and integrity, particularly concerning model poisoning and data integrity. Specifically, the cut layer selection significantly influences the system’s susceptibility to model poisoning attacks. For instance, extensive client-side training elevates the risk of client-side poisoning, complicating the defense against distributed attacks. Thus, sophisticated aggregation techniques are required to counter model poisoning risks effectively. An inappropriately chosen cut layer can also hinder data authenticity and integrity validation, posing challenges to maintaining the model’s training integrity. The selection of the cut layer, therefore, demands careful consideration, balancing between minimizing security risks and maximizing the efficiency of SFL.
IV-C Lightweight Design
@@
To enhance the efficiency of SFL, it is beneficial to consider the quantization approach. In the context of SFL, adopting a quantization approach can substantially reduce the size of both local model updates and intermediate data, which is useful particularly in resource-constrained clients. This reduction has several advantages, including reduced bandwidth usage, lower transmission energy, and reduced storage needs. Furthermore, quantization introduces inherent privacy guarantees. This is because reducing the precision of transmitted data inherently obfuscates the exact values, thereby adding an extra layer of protection for sensitive information, which is crucial in maintaining data privacy. Moreover, to mitigate the burden of training on resource-constrained clients with reduced energy consumption, the strategy of downsizing the resolution and the total volume of training data can be employed, considering different data distributions and acceptable accuracy levels. Additionally, selecting a lightweight DL model or employing model pruning techniques is adopted to facilitate a lightweight design for SFL.
V Conclusions